# Resonant spin Hall effect in a nano-ribbon of a spin-orbit coupled electron system

### v1

@@ Abstract
We present a theoretical study of spin Hall phenomena in a nanoribbon made of a two-dimensional square lattice with Rashba and Dresselhaus spin-orbit coupling. We show that the nanoribbon can give rise to a number of additional spin degeneracy points as well as anti-crossing points, apart from the $\Gamma$ point, between two nearest sub-bands. We compute the spin Hall conductivity and demonstrate that it diverges and gives rise to a resonance when chemical potential passes through those spin-degenerate or anti-crossing points. Contrary to the previous studies, here such resonance emerges even without any external perturbation like magnetic field or light. In addition, we examine the influence of anisotropy in Rashba and Dresselhaus interactions, as well as finite-temperature effects, and show that the inter subband resonance remains robust. Finally, we also investigate the signature of such additional spin degeneracy and anti-crossing points in the longitudinal conductance by using retarded Green function approach in lattice model. The peculiar features of the bands are reflected in the longitudinal conductance, which takes quantized values of $2ne^{2}/{h},$ where $n$ denotes the number of bands occupied by the chemical potential with each band having spin split sub-bands.
@@ I Introduction
The spin Hall effect (SHE)[1, 2, 3, 4] is the hall mark of spin-orbit coupling term in a two-dimensional (22D) fermionic system in which fermions with opposite spins are scattered to opposite transverse edge normal to the applied bias. The spin-orbit interaction is the result of relativistic effect, and in a 22D electronic system it mainly arises due to inversion symmetry breaking field across the semiconductor heterojunctions[5, 6]. There are two types of spin-orbit interactions, namely Rashba spin-orbit interaction (RSOI) and Dresselhaus spin-orbit interaction (DSOI). The RSOI arises due to the lack of structural inversion symmetry across the quantum well in a semiconductor heterojunction [7] , whereas the DSOI is the result of bulk inversion symmetry breaking [8]. The RSOI is externally tunable by applying a gate voltage [9, 10, 11], whereas the DSOI is an intrinsic property of the system that cannot be enhanced or tuned through external perturbations. The spin-orbit interaction removes the spin-degeneracy even in the absence of a magnetic field. Typical materials that exhibit RSOC include indium-based compounds such as InAs and heterostructures like GaInAs/GaAlAs. The degree of spin separation along the transverse direction to the applied bias is generally quantified by spin Hall conductivity (SHC) that was predicted to exhibit a universal constant value in 2D electron gas (22DEG) with $k$-linear RSOC[2]. Subsequently, the SHC in a 22D heavy-hole gas (22DHG) with $k$-cubic RSOI was also carried out and found that instead of universal constant value it’s sensitive to the strength of RSOI[14, 15].
@@
Apart from the above mentioned works on SHE, usual integer quantum Hall phenomena in such spin-orbit coupled system were also carried out [16, 17, 18, 19, 20, 21]. A series of theoretical investigations on spin-related transport phenomena in 22DHG in presence of weak magnetic field have also been studied[22, 23, 24, 25]. Moreover, the search for the materials or mechanisms to achieve giant or large SHC[26, 27, 28] continues to be an active topic in the field of spintronics.
@@
It is noteworthy to mention that the SHE is one of the key signatures of spin-Hall edge modes in the quantum-spin-Hall liquid discovered in 2D Dirac materials with spin-orbit coupling[29, 30]. The key difference with quantum spin-Hall insulator is that it belongs to $\mathbb{Z}_{2}$ topological insulator where bulk is insulating gapped with counter propagating spin-dependent gapless edge modes. Whereas in semiconductor heterostructures, the SHE is not associated with toplogy and its bulk is neither gapped nor it is due to spin-dependent edge modes. The SHE in semiconductor heterostructure is purely due to momentum dependent spin sensitive Lorentz force aided by spin-orbit interaction. From the experimental perspective, semiconductor compounds are still much more reliable platform than 2D Dirac materials to realize SHE.
@@
The present work is motivated by the phenomena called spin Hall resonance based on spin-orbit coupled quantum Hall system, predicted by Shen. et al. [31], in which the spin Hall conductivity exhibits a resonance if spin splitting vanishes between two nearest Landau levels with opposite spin at chemical potential. Such resonance was also predicted in a cubic Rashba system [32]. Recently, one of us has shown that even without Landau level such resonance can occur by applying linearly polarized light, in a $k$-cubic Rashba system [33]. In this work, our objective is to seek an answer of the question- is it possible to achieve such resonance even without any external perturbations like magnetic field or light? We show that such resonance can occur between two nearest sub-bands with opposite spin in a finite-size system where the width is relatively much smaller compared to length. In fact, we note that a large number of spin-splitting vanishing points and anti-crossing points can emerge at different energy ranges among the higher sub-bands. We discretize the system on a square lattice and consider both normal-edge and zigzag-edge nanoribbons. We analyze their band spectrum and subsequently study the spin Hall conductivity in both types of nanoribbon.
@@
This paper is organized as follows. In Sec. II, we introduce the model Hamiltonian and its discretization. Section III is devoted to the band structure of the nanoribbon. The Transport properties along with the effects of anisotropy and finite temperature on the spin Hall conductivity are presented and discussed in Sec. IV. and finally, Sec. V summarizes our findings and provides concluding remarks.
@@ II Model Hamiltonian and Discretization
In this section, we briefly discuss the electronic system first. Our objective is to study the spin transport with the finite size effects in a 22D electronic system. In order to include these effects, we need to work in a real space basis by discretizing the well-known spin-orbit coupled low energy effective Hamiltonian[7]-[8]
@@
where $\partial_{j}=\partial/\partial j$ ($j=x,y$), $\alpha$ and $\beta$ are the strength of the Rashba and the Dresselhaus spin-orbit interactions, $\bm{\sigma}=\{\sigma_{x},\sigma_{y}\}$ are Pauli matrices in real spin basis, and $m^{\ast}$ is the effective mass of the electron. In order to discretise the Hamiltonian, we shall assume a square lattice with the lattice periodicity $a$ along both the $x$- and $y$- directions. We use the plane wave representation of annihilation operator in the basis of lattice sites $(j,l)$, $c_{\bm{k},\sigma}\;=\;\frac{4\pi^{2}}{a^{2}}\sum_{j,l}e^{-ia(k_{x}j+k_{y}l)}\,c_{j,l,\sigma}$, we obtain the substitution rules for the continuum momentum operators (with $\lambda=x,y$)
@@
In the similar fashion, the RSOI and DSOI terms can also be represented in discrete lattice space. Using Eq. (2) in the Rashba and Dresselhaus interaction terms in Eq. (1), we obtain the second quantized form of the Rashba spin-orbit interaction [34] and Dresselhaus spin-orbit interaction [35] as
@@ III Energy Spectrum of the Nanoribbon
Having established the tight-binding Hamiltonian in real space and its discretization on a square lattice, we now turn to the nanoribbon geometry, which is the central focus of this work. In particular, we analyze two types of nanoribbons, namely straight-edge and zigzag-edge geometries as shown in Fig. 1(a) and Fig. 1(b), respectively. To obtain the energy spectrum for nanoribbons with both types of edge terminations, we construct a Hamiltonian analogous to the infinite lattice case [36, 37]. We define a nanoribbon as a stripe of finite width along the $x$-direction whereas infinitely extended along the $y$-direction. Because of the translational symmetry and lattice periodicity along the $y$-direction, we can write down the total Hamiltonian by using Bloch’s theorem as
@@
where $H_{00}$ is the on-site hopping matrix including the spin-orbit coupling of each supercell as highlighted in red colour in Fig. 1, and $H_{10}$ denotes the hoping matrix between two nearest supercells. The width of the ribbon is described by the number of atomic sites in each supercell. Note that because of the translational symmetry along the $y$-direction, $k_{y}$ is a conserved quantity and acts as a good quantum number. The energy eigen values are computed numerically by diagonalizing Eq. (10). The tight-binding Hamiltonian, including spin degrees of freedom, yields $2N$ energy sub-bands.
@@
We plot energy spectrum for zigzag nanoribbon in presence of RSOI and DSOI in Fig. (2). For better visualisation, the energy spectrum is shown only for a few lowest sub-bands around the $\Gamma$ point. The energy spectrum is referenced from a baseline value of $4t$, which corresponds to a constant energy shift applied to the full Hamiltonian. The straight edge band spectrum is given in the Appendix, as its qualitative features are essentially the same.
In the bulk, spin degeneracy is guaranteed at the $\Gamma$ point, and we find that this characteristic is preserved in the nanoribbon, with all sub-bands remaining spin-degenerate at $\Gamma$ point. In addition, several extra spin degeneracy points where the spin splitting vanishes emerge within the same sub-band. Moreover, number of spin degeneracy points as well as anti-crossing points appear between two nearest sub-bands with opposite spin at different $k_{y}$ points. It is important to note that anti-crossing points are distinct from spin degeneracy points, in the latter, the spin splitting vanishes, whereas in the former the two bands approach each other very closely but a small energy gap remains.
Figure 2: Energy spectrum of zigzag edge nanoribbon for $\alpha=0.1$ and $\beta=0.09$ in units of tata. The red and blue curves denote spin-up and spin-down branches. The anti-crossings and spin-degeneracy between two nearest sub-bands are shown by squares and circles, respectively.
We attribute these additional degeneracies and anticrossings to the competition between the RSOI and DSOI i.e., only Rashba or Dresselhaus alone cannot give rise to such degeneracy points and anti-crossing points. This is very much similar to the case of Rashba spin-orbit coupled quantum Hall system [31], where such additional spin-degeneracy is the result of the competition between the Rashba and Zeeman splitting. However, in the present case the Zeeman splitting is absent and its role is played by DSOI, whereas the Landau level index is played by finite width induced energy sub-bands. In zigzag-edge nanoribbon, increasing $\beta$ enhances the strength of inter-band anticrossings. In particular, for $\beta=0.09$ [Fig. 2], the spin-split sub-bands show visibly larger anticrossings, whereas for smaller $\beta$ the anticrossings remain relatively weak.
@@ IV Transport Properties of the Nanoribbon
IV.1 Spin Hall conductivity
The spin Hall conductivity is calculated by using linear response based Kubo formula as used in Ref.[2]. However, in the nanoribbon case, the system is infinite along the $y$ direction and preserves the translational symmetry, whereas the width is restricted along the $x$- direction breaking the translational symmetry. Hence, $k_{y}$ remains the conserved quantity and acts as a good quantum number . Hence an electronic state can be fully described by $\zeta=\{s,k_{y},n\}$ where $n$ is the sub-band index and s=+()s=+(-) denotes spin-up (down). The SHC can now be written for a nanoribbon as
@@ IV.1 Spin Hall conductivity
where $\Omega=L_{x}\times L_{y}$ denotes the area, $f_{\zeta}=[1+\exp\{(E_{\zeta}-\mu)/(k_{B}T)\}]^{-1}$ is the Fermi-Dirac distribution function with $\mu$ being the chemical potential. Also, $\hat{J}_{x}^{z}$ is the spin current operator and is given by $\hbar\{\sigma^{z},\hat{v_{x}}\}/4$, where the velocity operators are obtained $\hat{v}_{y}=(\hbar)^{-1}\partial H/\partial k_{y}$ and $\hat{v}_{x}=(i\hbar)^{-1}[H,x]$ with $x=\sum_{n}c_{n}^{\dagger}\,n\,c_{n}$ [38]. In the above Eq. 11 we perform $\sum_{\zeta\neq\zeta^{\prime}}\rightarrow(L_{y}/2\pi)\int dk_{y}\sum_{n,n^{\prime},s,s^{\prime}}$ to calculate the SHC for the nanoribbon geometry.
@@
Figure 3: Band structure (left) and spin Hall conductivity normalized with $\sigma_{0}=e/(8\pi)$ for a zigzag edge nanoribbon with width $N=21$ and $\alpha=0.1$, $\beta=0.09$ in units of tata at temperature $k_{B}T/t=10^{-5}$. The SHC exhibits resonance at various $\mu$. The brown dashed lines highlight the resonance due to interband anti-crossings, while the black dashed lines indicate the resonance due to vanishing spin splitting.
@@
We evaluate the SHC for zigzag nanoribbon and straight edge nanoribbons. The zigzag case is shown in the Fig. 3 whereas the straight edge is not shown here rather shifted to appendix A as qualitative nature remain same. We observe that the SHC is enhanced sharply at a number of chemical potential, which correspond to the inter-band spin splitting vanishing or anti-crossing points. This enhanced peaks are nothing but the so-called spin Hall resonance which precisely occur at the energies corresponding to the inter-band spin degeneracy points or anti-crossing points in the band structure. Importantly, we also check that in the limiting cases of a pure Rashba system ($\beta=0$) or a pure Dresselhaus system ($\alpha=0$), the inter-band spin degeneracy points are strongly suppressed or absent, and as a result, the SHC does not exhibit any resonant peaks. This observation suggests that a relative competition between two types of spin-splitting is required to achieve such resonance, here it is caused by RSOI and DSOI.
@@
Our findings resonate well with the theoretical predictions made by Shen et al. [31], where they studied resonant spin Hall conductance in 2DEGs under a perpendicular magnetic field. In that work, the RSOI competes with the Zeeman splitting to induce energy level crossings that give rise to the resonance. In our case, an analogous mechanism takes place in nanoribbons via the interplay between RSOI and DSOI, resulting in anti-crossings and degeneracy points between spin-resolved sub-bands and subsequent enhancement of the SHC.
@@
Figure 4: Spin Hall conductivity versus chemical potential for zigzag edge nanoribbon with width $N=21$ and $\alpha=0.1$, $\beta=0.09$ in units of tata for various temperatures.
@@
It is noteworthy to mention that when $\alpha=\beta$, the system exhibits a high degree of symmetry, and although many anti-crossings are visible in the band structure, the matrix elements of the spin current operator vanishes. Consequently, the SHC is identically zero across the entire range of chemical potentials. We also comment that as the spin Hall phenomena occurs in a $2D$ fermionic system, it is desirable to keep width also quite large compared to the present case. Here, we keep width very narrow as $N=21$ only to make degeneracy or anti-crossing points clearly visible. With the further increase of width a large number of sub-bands will be formed and many such degeneracy or anti-crossing points exhibiting resonance in SHC will occur.

### v2

# Resonant spin Hall and Nernst effect in a nanoribbon of a spin-orbit coupled electronic system

@@ Abstract
We present a theoretical study of spin Hall phenomenon in a nanoribbon of a two-dimensional electronic system with Rashba and Dresselhaus spin-orbit coupling. We model the electronic system by a square lattice in real space. We show that such nanoribbon can give rise to a number of additional spin degeneracy points as well as anticrossing points, apart from the $\Gamma$ point, between two opposite spin subbands. We compute the SHC and demonstrate that it diverges and gives rise to a resonance when the chemical potential passes through those spin degenerate or anticrossing points. Contrary to the previous studies, here such resonance emerges even without any external perturbation like magnetic field or light. We also examine the spin Nernst effect and find that it shows clear peaks at the anticrossing and spin degeneracy points, consistent with the Mott relation at low temperature. Finally, we also investigate the signature of such additional spin degeneracy and anticrossing points in the longitudinal conductance by using the retarded Green function approach in lattice model. The finite width induced subbands are reflected in the longitudinal conductance, which takes quantized values of $2ne^{2}/{h}$ where $n$ denotes the number of bands occupied by the chemical potential with each band having spin split subbands. We also note that anticrossing that occurs at low energy between two opposite spin subbands could be also detected via longitudinal conductance.
@@ I Introduction
The spin Hall effect (SHE) Hirsch (1999); Sinova et al. (2004, 2015); Murakami et al. (2003) is the hall mark of spin-orbit coupling (SOC) term in a two-dimensional (22D) fermionic system in which fermions with opposite spins are scattered to opposite transverse edge normal to the applied bias. The spin-orbit interaction is the result of relativistic effect, and in a 22D electronic system it mainly arises due to inversion symmetry breaking field across the semiconductor heterojunctions Winkler (2003); Žutić et al. (2004). There are two types of spin-orbit interactions, namely Rashba spin-orbit interaction (RSOI) and Dresselhaus spin-orbit interaction (DSOI). The RSOI arises due to the lack of structural inversion symmetry across the quantum well in a semiconductor heterojunction Bychkov and Rashba (1984) , whereas the DSOI is the result of bulk inversion symmetry breaking (Dresselhaus, 1955). The RSOI is externally tunable by applying a gate voltage Engels et al. (1997); Nitta et al. (1997); Heida et al. (1998), whereas the DSOI is an intrinsic property of the system that cannot be enhanced or tuned through external perturbations. The spin-orbit interaction removes the spin degeneracy even in the absence of a magnetic field. Typical materials that exhibit RSOI include indium-based compounds such as InAs and heterostructures like GaInAs/GaAlAs Scherübl et al. (2016); Nitta et al. (1997); Engels et al. (1997); Matsuyama et al. (2000). The degree of spin separation along the transverse direction to the applied bias is generally quantified by spin Hall conductivity that was predicted to exhibit a universal constant value in 2D electron gas (22DEG) with $k$-linear RSOI Sinova et al. (2004). Subsequently, the spin Hall conductivity in a 22D heavy-hole gas (22DHG) with $k$-cubic RSOI was also carried out and found that instead of universal constant value it’s sensitive to the strength of RSOI Murakami (2004); Schliemann and Loss (2005). Several experimental works also reported SHE phenomenon in semiconductor heterostructure  Kato et al. (2004); Wunderlich et al. (2005); Valenzuela and Tinkham (2006); Kimura et al. (2007).
@@
Apart from the above mentioned works on SHE, usual integer quantum Hall phenomenon in such spin-orbit coupled system were also carried out  Das et al. (1990); Wang and Vasilopoulos (2003); Mawrie et al. (2014, 2017); Faridi et al. (2018). In fact, the presence of both the RSOI and DSOI have been also considered in several theoretical works Ramaglia et al. (2006); Vasilopoulos and Wang (2006); Moca and Marinescu (2005); Wang and Vasilopoulos (2005); Lucignano et al. (2008); Farooq et al. (2022). A series of theoretical investigations on spin-related transport phenomena in 22DHG in presence of weak magnetic field have also been studied Liu et al. (2018); Marcellina et al. (2020, 2018); Cullen et al. (2020). Moreover, the search for the fermionic system or mechanisms to generate giant or large SHC Zhu et al. (2019); Yen and Guo (2020); Lou et al. (2020) continues to be an active topic in spintronics.
@@
It is to be noted that the SHE is also one of the key signatures of spin-Hall edge modes in the quantum-spin-Hall liquid discovered in 2D Dirac materials with spin-orbit coupling Kane and Mele (2005); Ezawa (2012). The key difference with quantum spin-Hall insulator is that it belongs to $\mathbb{Z}_{2}$ topological insulator where bulk is insulating gapped with counter propagating spin-dependent gapless edge modes, whereas semiconductor heterostructure is not topological insulator.
While the SHE describes the generation of transverse spin currents driven by an electric field, its thermal counterpart, The spin Nernst effect (SNE) arises when a longitudinal temperature gradient produces a transverse spin current through the same underlying SOC induced deflection mechanisms. In analogy with the spin Hall conductivity, the strength of this thermally driven response is characterized by the spin Nernst coefficient (SNC), which captures how efficiently a material converts thermal bias into transverse spin flow. Earlier works in mesoscopic crossbar with RSOI has demonstrated that a thermal gradient generates transverse spin current, and the spin Nernst response exhibits strong dependence on Fermi level, disorder, and external Magnetic field. In magnetic field driven systems, the Nernst type response was found to display sharp enhancements whenever the chemical potential crosses Landau levels Cheng et al. (2008). In graphene nanoribbon geometries, it has been found that the Nernst response strongly depends on ribbon width and edge chirality Xing et al. (2009). More recently, first principles Berry phase calculations in noncollinear antiferromagnets have predicted very large SNC values Guo and Wang (2017). Additionally, considerable effort is being devoted to an electronic system with RSOI in confined geometries to explore the interplay of quantum confinement and spin-degree of freedom in transport properties Maiellaro et al. (2025); Mondal et al. (2025). Hence electronic systems with RSOI continues to be an active research field, particularly in the context of spin dependent quantum phenomena Chen et al. (2025); Poshakinskiy (2025).
@@
The present work is motivated by the phenomenon called spin Hall resonance based on spin-orbit coupled quantum Hall system, predicted by Shen et al. Shen et al. (2004), in which the spin Hall conductance (SHC) exhibits a resonance if spin splitting vanishes between two nearest Landau levels with opposite spin at chemical potential. This study was also extended to a 22DEG with the $k^{3}$-RSOI Ma and Liu (2006), and in a 22DEG with the presence of both, the RSOI and DSOI Zhang et al. (2008). Recently, one of us has shown that even without Landau level such resonance can occur by applying linearly polarized light, in a $k$-cubic Rashba system  Bhattacharya and Islam (2021). We also note that anticrossings between nearest subbands with opposite spin polarization, accompanied by peaks in the SHC, were previously reported in electronic systems with RSOI in the presence of an additional fictitious lateral parabolic confinement potential Moroz and Barnes (1999). However, a parabolic confinement potential is also the result of perpendicular magnetic field. Hence, these two works conceptually resemble to each other, except in presence of magnetic field additional splitting (Zeeman splititng) competes with RSOI and yields a complete vanishing of spin splitting and resulting resonance in the SHC Shen et al. (2004). Whereas, in presence of a fictitious confinement potential there is no Zeeman term and resonance does not occur, rather several anticrossing appears. We show that such resonance can occur between two subbands with opposite spin in a finite-size system where the width is relatively much smaller compared to length. In fact, we note that a large number of spin splitting vanishing points and anticrossing points can emerge at different energy ranges among the higher subbands. In this work, we discretize the system on a square lattice and consider both normal edge and zigzag edge nanoribbons. We analyze their band spectrum and subsequently study the spin Hall conductivity and spin Nernst effect. We show that for a finite width with RSOI and DSOI, a number of crossing and anticrossing occurs between two subbands with opposite spins, yielding resonance in SHC and Nernst effect.
@@
This paper is organized as follows. In Sec. II, we introduce the model Hamiltonian and its discretization. Section III is devoted to the band structure of the nanoribbon. The Transport properties along with the effects of anisotropy and finite temperature on the SHC are presented and discussed in Sec. IV. and finally, Sec. V summarizes our findings and provides concluding remarks.
@@ II Model Hamiltonian and Discretization
In this section, we briefly discuss the electronic system first. Our objective is to study the spin transport with the finite size effects in a 22D electronic system. In order to include these effects, we need to work in a real space basis by discretizing the well-known spin-orbit coupled low energy effective Hamiltonian Bychkov and Rashba (1984); Dresselhaus (1955)
@@
where $\partial_{j}=\partial/\partial j$ ($j=x,y$), $\alpha$ and $\beta$ are the strength of RSOI and DSOI respectively, $\bm{\sigma}=\{\sigma_{x},\sigma_{y}\}$ are Pauli matrices in real spin basis, and $m^{\ast}$ is the effective mass of the electron. A number of works have already considered lattice model of the above Hamiltonian Moca and Marinescu (2005); Tanaka and Sato (2024). However, we briefly outline the discretization of the above mentioned Hamiltonian by considering a square lattice with lattice constant a along both the $x$- and $y$- directions. We use the plane wave representation of annihilation operator in the basis of lattice sites $(j,l)$, $c_{\bm{k},\sigma}\;=\;\frac{4\pi^{2}}{a^{2}}\sum_{j,l}e^{-ia(k_{x}j+k_{y}l)}\,c_{j,l,\sigma}$, we obtain the substitution rules for the continuum momentum operators (with $\lambda=x,y$)
@@
In the similar fashion, the RSOI and DSOI terms can also be represented in discrete lattice space. Using Eq. (2) and Eq. (1), we obtain the second quantized form of the RSOI Mireles and Kirczenow (2001) and DSOI Moca and Marinescu (2005) as
@@ III Energy Spectrum of the Nanoribbon
Having established the tight-binding Hamiltonian in real space and its discretization on a square lattice, we now turn to the nanoribbon geometry, which is the central focus of this work. In particular, we analyze two types of nanoribbons, namely straight edge and zigzag edge geometries as shown in Fig. 1(a) and Fig. 1(b), respectively. To obtain the energy spectrum for nanoribbons with both types of edge terminations, we construct a Hamiltonian analogous to the infinite lattice case Islam et al. (2018); Dutta et al. (2013). We define a nanoribbon as a stripe of finite width along the $x$-direction whereas infinitely extended along the $y$-direction. Because of the translational symmetry and lattice periodicity along the $y$-direction, we can write down the total Hamiltonian by using Bloch’s theorem as
@@
where $H_{00}$ is the on-site hopping matrix including the spin-orbit coupling of each supercell as highlighted in red colour in Fig. 1, and $H_{10}$ denotes the hoping matrix between two nearest supercells. The width of the ribbon is described by the number of atomic sites in each supercell. Note that because of the translational symmetry along the $y$-direction, $k_{y}$ is a conserved quantity and acts as a good quantum number. The energy eigen values are computed numerically by diagonalizing Eq. (10). The tight-binding Hamiltonian with spin degrees of freedom gives rise to $N$ transverse bands. In the presence of RSOI and DSOI, each band is split into two in-plane spin-polarized subbands, leading to a total of $2N$ energy subbands. By denoting each subband by $n$, an energy eigen ket can be fully described by $|\zeta\rangle=\{s,k_{y},n\}$ where s=+()s=+(-) denotes spin-up (down) subband, respectively.
@@
We plot spin resolved energy spectrum for zigzag nanoribbon around $\Gamma$ point in Fig. 2 for lowest few subbands. The in-plane spin polarization is defined as $\langle\mathbf{S}\rangle=(\hbar/2)\langle\zeta|\bm{\sigma}|\zeta\rangle$. In this work, the terms spin up and spin down refer to states with opposite projections of $\langle\bm{\sigma}\rangle$ which are distinguished by different colours in the band structure. The energy spectrum is referenced from a baseline value of $4t$, which corresponds to a constant energy shift applied to the full Hamiltonian. The straight edge band spectrum is given in the Appendix A, as it’s qualitative features are essentially the same for lowest few bands.
In the bulk system, spin degeneracy is guaranteed at the $\Gamma$ point by time reversal symmetry (TRS). This characteristic is preserved in the nanoribbon geometry because the confinement does not break TRS. Therefore, TRS imposes Kramers degeneracy $E_{n,k_{y},\uparrow}=E_{n,-k_{y},\downarrow}$ so that all subbands remain spin degenerate at $k_{y}=0$. In addition, several extra spin degeneracy points, at which the spin splitting vanishes, emerge between different subbands with opposite spin. Furthermore, multiple anticrossing points between subbands with opposite spin polarization appear as well. It is important to emphasize that anticrossing points are distinct from spin degeneracy points. At a spin degeneracy point the spin splitting vanishes, yielding a degeneracy of the same order as that observed at $k_{y}=0$. In contrast, at an anticrossing point the two bands approach each other closely but remain separated by a finite energy gap, indicating an avoided crossing with a clearly resolvable splitting.
Figure 2: Spin resolved band dispersion is plotted for a zigzag edge nanoribbon with $\alpha=0.1$ and $\beta=0.09$ in units of tata. The colour represents the expectation value of the in-plane spin operator, i.e., the spin polarization in units of $\hbar/2$. Anticrossing points are highlighted by squares, while spin degeneracy points are marked by circles.
We attribute these additional degeneracies and anticrossings to the competition between the RSOI and DSOI i.e., only RSOI or DSOI alone cannot give rise to such degeneracy points and anticrossing points. This is very much similar to the case of Rashba spin-orbit coupled quantum Hall system Shen et al. (2004), where such additional spin degeneracy is the result of the competition between the Rashba and Zeeman splitting. However, in the present case the Zeeman splitting is absent and its role is played by DSOI, whereas the Landau level index is played by finite width induced energy subbands. In zigzag edge nanoribbon, increasing $\beta$ enhances the strength of inter-band anticrossings. In particular, for $\beta/ta=0.09$ [Fig. 2], the spin-split subbands show visibly larger anticrossings, whereas for smaller $\beta$ the anticrossings remain relatively weak.
@@ IV Transport Properties of the Nanoribbon
IV.1 Spin Hall conductance
The SHC is calculated by using linear response based Kubo formula as used in Ref. Sinova et al. (2004). The SHC can now be written for a nanoribbon as
@@ IV.1 Spin Hall conductance
where $\Omega=L_{x}\times L_{y}$ denotes the area, $f_{\zeta}=[1+\exp\{(E_{\zeta}-\mu)/(k_{B}T)\}]^{-1}$ is the Fermi-Dirac distribution function with $\mu$ being the chemical potential. Also, $\hat{J}_{x}^{z}$ is the spin current operator and is given by $\hbar\{\sigma^{z},\hat{v_{x}}\}/4$, where the velocity operators are obtained $\hat{v}_{y}=(\hbar)^{-1}\partial H/\partial k_{y}$ and $\hat{v}_{x}=(i\hbar)^{-1}[H,x]$ with $x=\sum_{n}c_{n}^{\dagger}\,n\,c_{n}$ Maiti (2011). In the above Eq. (11) we perform $\sum_{\zeta\neq\zeta^{\prime}}\rightarrow(L_{y}/2\pi)\int dk_{y}\sum_{n,n^{\prime},s,s^{\prime}}$ to calculate the SHC for the nanoribbon geometry.
@@
Figure 3: Spin resolved band structure (left) and SHC (normalized by $\sigma_{0}=e/8\pi$) are shown for a zigzag edge nanoribbon with width $N=21$ and $\alpha=0.1$, $\beta=0.09$ in units of tata at temperature $k_{B}T/t=10^{-5}$. The spin Hall resonances arising from anticrossings occur at chemical potentials $\mu/t=0.115,0.3,$ and $0.58$, highlighted by brown dashed lines. The resonance associated with vanishing spin splitting appears at $\mu/t=0.65$, indicated by a black dashed line.
@@
We evaluate the SHC for zigzag edge nanoribbon and straight edge nanoribbon. The zigzag edge case is shown in the Fig. 3 whereas the straight edge is not shown here rather shifted to appendix A as qualitative nature remain same. We observe that the SHC is enhanced sharply at a number of chemical potential, which correspond to the inter-band spin splitting vanishing or anticrossing points. This enhanced peaks are nothing but the so-called spin Hall resonance which precisely occur at the energies corresponding to the inter-band spin degeneracy points or anticrossing points in the band structure. Next, we also check that in the limiting cases of a pure Rashba system ($\beta=0$) or a pure Dresselhaus system ($\alpha=0$), the inter-band spin degeneracy points are strongly suppressed or absent, and as a result, the SHC does not exhibit any resonant peaks. This observation suggests that a relative competition between two types of spin splitting is required to achieve such resonance, here it is caused by RSOI and DSOI. We also mention that, the spin Hall resonances persist for both even and odd lattice terminations. Although the boundary condition modifies the subband dispersions, the qualitative structure of the resonances remains the same.
@@
Our findings resonate well with the theoretical predictions made by Shen et al. (Shen et al., 2004), where they studied resonant SHC in 2DEGs under a perpendicular magnetic field. In that work, the RSOI competes with the Zeeman splitting to induce energy level crossings that give rise to the resonance. In our case, an analogous mechanism takes place in nanoribbons via the interplay between RSOI and DSOI, resulting in anticrossings and degeneracy points between spin-resolved subbands and subsequent enhancement of the SHC.
@@
Figure 4: Spin Hall conductance versus chemical potential is plotted for zigzag edge nanoribbon with width $N=21$ and $\alpha=0.1$, $\beta=0.09$ in units of tata for various temperatures.
We now comment on the physical origin of the different resonance widths visible in Fig. 3. The width of the spin Hall resonances is mainly determined by thermal broadening and by the extent of mixing between the participating subbands. When the mixing between subbands of different spin character extends over a wider interval of $k_{y}$, the contributing states are distributed over a larger momentum range, resulting in an increased resonance width. In contrast, when the mixing is confined to a narrow region of $k_{y}$, the contributing states are limited and the corresponding resonance remains sharp. This behavior is directly reflected in the Fig. 38, where broader resonances are observed at crossings and anticrossings characterized by an extended momentum range of spin mixing.
@@
It is noteworthy to mention that when $\alpha=\beta$, the system exhibits a high degree of symmetry, and although many spin degeneracy points are visible in the band structure of both zigzag and straight edge as shown in Fig. 11-13, the matrix elements of the spin current operator vanishes. Consequently, the SHC is identically zero across the entire range of chemical potentials. Also in such case the Hamiltonian commutes with the spin operator $\sigma_{x}-\sigma_{y}$ Bernevig et al. (2006), leading to the conservation of this spin component, this conserved spin symmetry forbids any net spin Hall response. This vanishing of SHC for $\alpha=\beta$ is not a finite size effect of nanoribbon but persists in the bulk system also Chang (2005). We also comment that, as the spin Hall phenomenon occurs in a 22D fermionic system, it is desirable to keep width also quite large compared to the present case. Here, we keep width very narrow as $N=21$ only to make degeneracy or anticrossing points clearly visible. With the further increase of width a large number of subbands will be formed and many such degeneracy or anticrossing points exhibiting resonance in SHC will occur.